# Toward Human-Quantum Computer Interaction: Interface Techniques for Usable Quantum Computing

### v2

@@ 6.3. Techniques for Machine Selection
(T2c) Exporting reusable code. The online platforms of QPU providers offer visualizations for machine properties, but developers need to make API calls to use them for their work, which is a tedious task as discussed earlier. Thus, our QPU data dashboard generates a reusable code snippet for selected properties of a QPU (P3: Baseline usability support). A code snippet takes a reusable function format so as to ensure its reusability in obtaining the same information for a different QPU.
@@ 6.4. Techniques for Circuit Review and Optimization
(T3a) Linking logical and physical circuits. QC developers who are testing or learning optimization strategies need to be able to inspect details (P2: Support at different scales). Thus, our interface juxtaposes a logical circuit and the optimized physical circuit. Furthermore, to support directly comparing the logical and physical circuits (P1: Link concept and quantum), the interface enables cross-highlighting of corresponding logical and physical gates on a circuit diagram. We note that there is no deterministic way to map logical and physical circuits, and the most exhaustive method (brute-force) that compares every $x$-consecutive gates has factorial (!) time complexity while not guaranteeing a perfect match. Thus, we applied a greedy approach that compares the first $k$ gates to the first $k-1$ gates, reducing the time complexity to a polynomial time. Our Supplementary Material further details this.
@@
(T3d) Showing fidelity information. A benefit of mapping gates on a machine diagram is directly showing how the fidelity of a circuit on a QPU decays over time. Fidelity information means how gate error accumulates on qubits over time. For instance, it is best if accumulated gate error is well distributed across different qubits rather than concentrated on a few qubits. An approach to assessing such fidelity is estimated success probability (ESP) (nishio2020:esp), which is defined as the product of the success rates ($1-\text{error}$) of all the gates involved in a physical circuit. Thus, we integrated three visualizations within the above techniques: (1) the ESP value of each layer, (2) the cumulative ESP value at each layer, and (3) the cumulative ESP value of each qubit at each layer. The ESP value of layer $i$ refers to the ESP value for the gates and qubits in the layer $i$ only. The cumulative ESP value at layer $i$ is defined as the ESP value for the gates and qubits involved in the first $i$ layers. The cumulative ESP value of qubit $j$ at layer $i$ means the ESP value for the gates applied to qubit $j$ for the first $i$ layers. The first two methods appear along with a traditional circuit diagram (T3a), and the last method is shown on a machine diagram (T3b). Together with summary ESP values in the interface, these methods allow for inspecting ESP at diverse levels (P2: Support at different scales).
(T3e) Comparing optimizations. Because different factors of a quantum machine affects optimization outcomes (i.e., physical circuits), it is crucial for more than intermediate-level QC developers to be able compare different optimization. Thus, our circuit optimization interface allows for comparing multiple optimization outcomes of the same logical circuit. First, a table summarizes all the optimization results in terms of the number of gates, ESP values, etc. (P2: Support at different scales). Next, the interface allows for browsing detailed information of each optimization outcome like circuit diagrams and detailed fidelity information. When a gate in the logical circuit is selected, corresponding physical gates are highlighted across physical circuits, which is also summarized in a widget for a concise look.
@@ 6.5. Techniques for Result Analysis
(T4c) Hypothetical error adjustment. Given that QC program outputs include some noises occurring due to gate errors, the counts of these bit strings have uncertainty to a degree. Yet, checking where gate errors actually happened is not realistic and practically impossible because they happen within a few nanoseconds. When uncertainty is not easily tractable, a simulation-based method like Monte-Carlo approximation (mooney1997:monte) is widely used in statistics (christos2001:monteCarlo). Thus, we designed a hypothetical error adjustment method for approximating gate errors using Monte-Carlo simulation to directly present noise from a machine in a more interpretable form (P1: Link concept and quantum) and hence to provide better contexts for collaborative QC programming like benchmarking (P3: Baseline usability support). We provide detailed computation in Supplementary Material.
This simulation results in an uncertainty visualization like Figure 5. The orange ticks represent the measured count and the blue ticks indicate the mean of hypothetical error-adjusted counts. The black line behind a blue tick shows the 95% confidence interval of the corresponding hypothetical error-adjusted count. Note that this procedure has a general tendency to level the count values toward the center (i.e., the number of shots divided by the number of possible state vectors) because state vectors with more counts will decrease while those with less counts will increase. This visualization can help QC developers to assess whether they have ran the circuit with sufficient shots to differentiate state vector counts and if their circuit is returning what they have intended. For instance, Figure 5A shows a well-differentiated outcome while Figure 5B and C indicate the needs for updating a circuit or increasing the number of shots to obtain a clear outcome.
@@
(T4e) Export and retrieve job data. A job is a common term to refer to the execution and results data of a circuit. Sharing job data is highly important for collaborative QC development and benchmark. To share job data, a developer needs to manually combine the logical and physical circuit designs in a QASM format, state vector counts, and physical properties of a QPU used for the circuit. Thus, we developed APIs that combine those into a single data object; export into a shareable file; and retrieve the data from those file (P3: Baseline usability support). For example, once loaded, the job data object provides methods that easily visualize the QC outcomes, circuits, and machine properties (e.g., opening an interface for result analysis).
@@ 6.6. Implementation Details
We implemented the above interaction techniques as high-fidelity prototypes. We implemented a circuit writer for circuit composition techniques, a machine explorer for machine selection techniques, a circuit viewer for circuit review and optimization techniques, and a result viewer for result analysis techniques. We built these interfaces using Svelte.js (svelte) and embedded them in Jupyter Notebook (jupyterNotebooks) using AnyWidget (anywidget). These interfaces are available as open-sourced libraries with demo notebooks [^7] https://github.com/see-mike-out/patoka or pip install patoka..
@@ 7.1. Case 1: Learning Shor’s Algorithm
First, James imports an circuit writer for an AerSimulator, a sandbox simulator with an unlimited number of qubits offered by Qiskit (Figure 6A1). In the interface, James sets automated qubit selection (T1c) given that he is not yet aware of how many qubits the circuit will need. After selecting ‘Shor’s algorithm’ subroutine, James sets a base factor to 2 and a divider to 15 (T1a: Problem-oriented circuit writing). As James updates the inputs for this subroutine, the interface verifies the circuit (T1b) and tells whether James provided required inputs and they are coherent. With confidence, thus, James extracts the actual Qiskit code from the interface (Figure 6A1), which he can paste into the next cell in the current notebook. The code includes all the required packages, comments for descriptions, circuit design, and optimization and execution functions.
@@
Running a circuit on a cloud QPU and obtaining the job data takes some time (up to a few days). Plus, reformatting the loaded data from the cloud is cumbersome. Thus, James calls a function (processJobData in Figure 6C1) to save the job once it is done, and retrieves it (retrieveJobData in Figure 6C2; T5e). Then, James loads a result viewer that includes meta data, traditional histogram, and natural number-based histogram. The meta information concisely tells him whether the circuit was run successfully (T5d) without looking for codes to retrieve the data (T0a, T0b). The histogram with natural numbers (Figure 6C3; T5b: problem-specific visualization) helps James to easily figure out the potential answers, 48 (mod $15=1$), 176 (mod $15=1$), and so on. Given the large output data size (39.4MB), Qiskit’s online platform on the browser failed to show the job outcome, whereas this interface can smoothly show it (T5a: Scalable data transmission). Lastly, to make sure that the machine was in a good status, James opens a machine explorer by passing a list of dates after and before the job. The machine explorer provides a pop-up to show the historical QPU data (T2b), as illustrated in Figure 7D.
@@ 7.2. Case 2: Running Quantum Machine Learning
An important topic in QML is developing subroutines, or ansatzs, that can be easily chained to achieve a larger ML network (ansatz). Paula is currently testing which filter to apply for her QML project. To do so, Paula opens an interactive circuit writer on her Jupyter notebook (Figure 8A). In this interface, Paula can provide the image file, the image color dimension, and filter type (T1a: Problem-specific circuit writing). This generates a code snippet for running this circuit, which saves few hundred lines of code if she did manually.
@@ 7.3. Case 3: Studying Circuit Optimization Strategies
Laura is interested in four optimization strategies, including a baseline case, because the original research relating to those strategies (Li et al. (li2019:sabre)) did not look at the Toffoli gate. After passing the optimization strategies as Qiskit code (Figure 10C1), Laura imports the circuit viewer (Figure 10C2). Laura compares duration, gate count, layers and cumulative ESP in the summary table (T3d: fidelity information & T3f: comparing optimizations), and then checks the actual results via cross-highlighting (T3a: linking logical & physical circuits). By doing so, Laura can confirm that the stronger optimization improves the feasibility, showing that this circuit is not a trivial case.
@@ 8.1. HCI + PL for QC
Programming language (PL) research has made programming more robust with type systems, logical with better semantics, and efficient with compilers. In doing so, HCI and PL can have a synergy as it has shown in classical computing. As reviewed by Chasins et al. (chasins2021:plhci), by learning from each other, PL grammars can better capture targeted tasks, and user interfaces can provide more compositional and principled interactions. Currently, we do not write a QC program using some specialized forms (e.g., punch cards) but by building formal expressions based on classical PLs (e.g., Python libraries). Yet, QC programming expressions currently lack a robust type system and well-defined semantics to encode various problems. For example, a set of qubits can primarily encode states or phases in a probabilistic way, and their outputs can be interpreted as top-k candidates or as an empirical distribution. A program can repeat a subroutine to iteratively update qubit states or to use the information multiple times (so as to control other qubits). Present QC tools do not offer methods to explicitly indicate those behaviors, but programmers need to come up with other tools like NumPy on their own.
Therefore, a robust PL system for QC program should capture how QC programmers conceptualize such behaviors. Research in PL for QC can benefit from HCI methods, such as analyzing patterns in programming (e.g., (denny2012:syntaxerror; whalley2021:debugging)), demonstrating them using diverse use cases (e.g., (satyanarayan:vega2016; kim:erie2023)), and operating user studies (e.g., (chen2024:cse; myers2004:npl; myers2016:hcpl)). Such future work will make our circuit writer and result viewer interfaces operate more robustly and efficiently as well as better match to conceptual ideas.
@@ 8.2. Standardizing Interoperable Representations
Consistent representation methods of data are also necessary for building interfaces, but missed currently in QC tools. For example, authentication services like ‘Sign in via Google’ can work via standardized data structures that concretely connect layers of data bases, security services, and interfaces. While implementing the prototypes, however, we frequently experienced inconsistency in API designs and their outcome formats. A major effort in our prototyping was formatting QC data in a consistent structure so that enable on-demand user interactions on data-heavy interfaces like the circuit viewer and machine explorer. Otherwise, ad-hoc conversion had to occur every time a user clicks an element, which would add up browser-side latency.
Future research in data representation methods can work at the inter-operation of two layers: deeper, storage-efficient methods and shallower, interface-efficient methods. First, deeper representations can support storing QC data in classical databases, efficiently sharing them among users, and minimizing latency in between-application communication. Next, shallow expressions can help interfaces developers be less concerned with backend-side data structures, reducing the needs for ad-hoc methods. Lastly, a suite of compilers will play a key role in making those representation methods operate with each other. For instance, data visualization designs are often expressed as concise design specs (deep), which are compiled to systematically expressed visual entities (shallow) that people see on screen. Standardizing QC representation methods at varying levels will facilitate the development of useful APIs and collaborative interfaces.
@@ 8.3. Domain-specific QC Tools
QC is a specialized computing resource, intending for domain experts to solve particular types of problems. While we prototyped multiple interfaces that included problem-oriented circuit writing and result visualization, an interesting direction can be one-stop shop interfaces for specific domain areas that select and combine different techniques. For example, a QC interface for drug discovery and chemistry could help researchers to easily set superposition for molecules, encode chemical reactions as an oracle, and show outcomes in a suitable format (e.g., expressing state vectors as formulas). Similarly, simulation problems for physics can offer an editor for physical systems (e.g., a set of ions, photons) and show outcomes as animations for how those systems develop over time. The scalability of those tools will be an important concern.
Future work in domain-specific QC tools may also need to look at the interplay between graphical and character user interfaces. While graphical interfaces can reduce the gap between computational artifacts and conceptual ideas, it can be cumbersome to manipulate a lot of object by hand (e.g., dragging a few hundred gates to encode a molecule). Instead, specifying a problem as abstract formulae can make user interaction easier than graphical manipulation. Future work could also consider the interplay between QC programming and representation methods. For instance, QC programming tools can allow for directly typing in chemical formulas, and representations methods can include data about timed evolution of a physical system.
@@ 8.4. Feasible Evaluation Methods
While our goal was to propose relatively generic techniques to provide blueprints for next-generation QC tools, each interface in our work would need a designated user study for further evaluation. Prior work has done user studies on highly narrow tasks (e.g., understanding a circuit (ruan2024:quantumeyes; wen2024:quantivine), interpreting quantum states (ruan2023:venus), reading error rates (ruan2023:vacsen)) with expert users who have done research in QC. Yet, future QC interfaces should be able to accommodate those who have limited backgrounds in QC but want to explore the space for potential interdisciplinary work. Evaluation studies then should be able to address the lack of sufficient QC knowledge in a feasible way; for instance, an hour-long lecture for a single-shot study is unlikely to work. Example approaches can include a single-shot user study with a well-scoped tutorial (e.g., modularization), observing how communities around tools (e.g., Qiskit user group) change in a long term, and deploying tools as course materials.

### v3

@@ 6.3. Techniques for Machine Selection
(T2c) Exporting reusable code. The online platforms of QPU providers offer visualizations for machine properties, but developers need to make API calls to use them for their work, which is a tedious task as discussed earlier. Thus, our QPU data dashboard generates a reusable code snippet for selected properties of a QPU (P3: Baseline usability support). A code snippet takes a function format to ensure its reusability for different QPUs.
@@ 6.4. Techniques for Circuit Review and Optimization
(T3a) Linking logical and physical circuits. QC developers who are testing or learning optimization strategies need to be able to inspect details (P2: Support at different scales). Thus, our interface juxtaposes a logical circuit and its optimized physical circuits. Furthermore, to support directly comparing logical and physical circuits (P1: Link concept and quantum), the interface enables cross-highlighting of corresponding logical and physical gates on a circuit diagram. We note that there is no deterministic way to map logical and physical circuits, and the most exhaustive method (brute-force) that compares every $x$-consecutive gates has factorial (!) time complexity while not guaranteeing a perfect match. Thus, we applied a greedy approach that compares the first $k$ gates to the first $k-1$ gates, reducing the time complexity to a polynomial time. Our Supplementary Material further details this.
@@
(T3d) Showing fidelity information. Mapping gates on a machine diagram makes it more explicit to display how the fidelity of a circuit on a QPU decays over time. Fidelity information means how gate error accumulates on qubits over time. For instance, it is best if accumulated gate error is well distributed across different qubits rather than concentrated on a few qubits. An approach to assessing such fidelity is estimated success probability (ESP) (Nishio et al., 2020), which is defined as the product of the success rates ($1-\text{error}$) of all the gates involved in a physical circuit. Thus, we integrated three visualizations within the above techniques: (1) the ESP value of each layer, (2) the cumulative ESP value at each layer, and (3) the cumulative ESP value of each qubit at each layer. The ESP value of layer $i$ refers to the ESP value for the gates and qubits in the layer $i$ only. The cumulative ESP value at layer $i$ is defined as the ESP value for the gates and qubits involved in the first $i$ layers. The cumulative ESP value of qubit $j$ at layer $i$ means the ESP value for the gates applied to qubit $j$ for the first $i$ layers. The first two methods appear along with a traditional circuit diagram (T3a), and the last method is shown on a machine diagram (T3b). Together with summary ESP values in the interface, these methods allow for inspecting ESP at diverse levels (P2: Support at different scales).
(T3e) Comparing optimizations. Because different factors of a quantum machine affects optimization outcomes (i.e., physical circuits), it is crucial for more than intermediate-level QC developers to be able to compare different optimization results. Thus, our circuit optimization interface allows for comparing multiple optimization outcomes of the same logical circuit. First, a table summarizes all the optimization results in terms of the number of gates, ESP values, etc. (P2: Support at different scales). Next, the interface allows for browsing detailed information of each optimization outcome like circuit diagrams and detailed fidelity information. When a gate in the logical circuit is selected, corresponding physical gates are highlighted across physical circuits, which is also summarized in a widget for a concise look.
@@ 6.5. Techniques for Result Analysis
(T4c) Hypothetical error adjustment. Given that QC program outputs include some noises occurring due to gate errors, the counts of these bit strings have uncertainty to a degree. Yet, checking where gate errors actually happened is not realistic and practically impossible because they happen within a few nanoseconds. When uncertainty is not easily tractable, a simulation-based method like Monte-Carlo approximation (Mooney, 1997) is widely used in statistics (Papadopoulos and Yeung, 2001). Thus, we designed a hypothetical error adjustment method for approximating gate errors using Monte-Carlo simulation to directly present noise from a machine in a more interpretable form (P1: Link concept and quantum) and hence to provide better contexts for collaborative QC programming like benchmarking (P3: Baseline usability support). We provide detailed computation in Supplementary Material.
This simulation results in an uncertainty visualization like Figure 5. The orange ticks represent the measured count and the blue ticks indicate the mean of hypothetical error-adjusted counts. The black line behind a blue tick shows the 95% confidence interval of the corresponding hypothetical error-adjusted count. Note that this procedure has a general tendency to level the count values toward the center (i.e., the number of shots divided by the number of possible state vectors) because higher state vector counts will decrease whereas lower counts will increase. This visualization can help QC developers to assess whether they have ran the circuit with sufficient shots to differentiate state vector counts and if their circuit is returning what they have intended. For instance, Figure 5A shows a well-differentiated outcome while Figure 5B and C indicate the needs for updating a circuit or increasing the number of shots to obtain a clear outcome.
@@
(T4e) Export and retrieve job data. A job is a common term to refer to the execution and results data of a circuit. Sharing job data is highly important for collaborative QC development and benchmark. To share job data, a developer needs to manually combine the logical and physical circuit designs in a QASM format, state vector counts, and physical properties of a QPU used for the circuit. Thus, we developed APIs that combine those into a single data object; export into a shareable file; and retrieve the data from that file (P3: Baseline usability support). For example, once loaded, the job data object provides methods that easily visualize the QC outcomes, circuits, and machine properties (e.g., opening an interface for result analysis).
@@ 6.6. Implementation Details
We implemented the above interaction techniques as high-fidelity prototypes. We implemented a circuit writer for circuit composition techniques, a machine explorer for machine selection techniques, a circuit viewer for circuit review and optimization techniques, and a result viewer for result analysis techniques. We built these interfaces using Svelte.js (sve, 2024) and embedded them in Jupyter Notebook (Kluyver et al., 2016) using AnyWidget (any, 2023). These interfaces are available as open-sourced libraries with demo notebooks [^7] https://see-mike-out.github.io/patoka or pip install patoka..
@@ 7.1. Case 1: Learning Shor’s Algorithm
First, James imports an circuit writer for an AerSimulator, a sandbox simulator with an unlimited number of qubits offered by Qiskit (Figure 6A1). In the interface, James sets automated qubit selection (T1c) given that he is not yet aware of how many qubits the circuit will need. After selecting ‘Shor’s algorithm’ subroutine, James sets a base factor to 7 and a divider to 15 (T1a: Problem-oriented circuit writing). As James updates the inputs for this subroutine, the interface verifies the circuit (T1b) and tells whether James has provided required inputs and they are coherent. With confidence, thus, James extracts the actual Qiskit code from the interface (Figure 6A1), which he can paste into the next cell in the current notebook. The code includes all the required packages, comments for descriptions, circuit design, and optimization and execution functions.
@@
Running a circuit on a cloud QPU and obtaining the job data takes some time (up to a few days). Plus, reformatting the loaded data from the cloud is cumbersome. Thus, James calls a function (processJobData in Figure 6C1) to save the job once it is done, and retrieves it (retrieveJobData in Figure 6C2; T5e). Then, James loads a result viewer that includes meta data, traditional histogram, and natural number-based histogram. The meta information concisely tells him whether the circuit was run successfully (T5d) without looking for codes to retrieve the data (T0a, T0b). The histogram with natural numbers (Figure 6C3; T5b: problem-specific visualization) helps James to easily figure out the potential answers, $7^{8}$ (mod $15=1$), $7^{12}$ (mod $15=1$), and so on. Given the large output data size (39.4MB), Qiskit’s online platform on the browser failed to show the job outcome, whereas this interface can smoothly show it (T5a: Scalable data transmission). Lastly, to make sure that the machine was in a good status, James opens a machine explorer by passing a list of dates after and before the job. The machine explorer provides a pop-up to show the historical QPU data (T2b), as illustrated in Figure 7D.
@@ 7.2. Case 2: Running Quantum Machine Learning
An important topic in QML is developing subroutines, or ansatzs, that can be easily chained to achieve a larger ML network (PennyLane, [n. d.]). Paula is currently testing which filter to apply for her QML project. To do so, Paula opens an interactive circuit writer on her Jupyter notebook (Figure 8A). In this interface, Paula can provide the image file, the image color dimension, and filter type (T1a: Problem-specific circuit writing). This generates a code snippet for running this circuit, which saves few hundred lines of code if she did manually.
@@ 7.3. Case 3: Studying Circuit Optimization Strategies
Laura is interested in four optimization strategies, including a baseline case, because the original research relating to those strategies (Li et al. (Li et al., 2019)) did not look at the Toffoli gate. After passing the optimization strategies as Qiskit code (Figure 10C1), Laura imports the circuit viewer (Figure 10C2). Laura compares duration, gate count, layers and cumulative ESP in the summary table (T3d: fidelity information & T3f: comparing optimizations), and then checks the actual results via cross-highlighting (T3a: linking logical & physical circuits). By doing so, Laura can confirm that the stronger optimization improves the feasibility, showing that this circuit is not a trivial case.
@@ 8.1. HCI + PL for QC
Programming language (PL) research has made programming more robust with type systems, logical with better semantics, and efficient with compilers. In doing so, HCI and PL can have a synergy as it has shown in classical computing. As reviewed by Chasins et al. (Chasins et al., 2021), by learning from each other, PL grammars can better capture targeted tasks, and user interfaces can provide more compositional and principled interactions. Currently, we do not write a QC program using some specialized forms (e.g., punch cards) but by building formal expressions based on classical PLs (e.g., Python libraries). Yet, QC programming expressions currently lack a robust type system and well-defined semantics to encode various problems. For example, a set of qubits can primarily encode states or phases in a probabilistic way, and their outputs can be interpreted as top-k candidates or as an empirical distribution. A program can repeat a subroutine to iteratively update qubit states or to use the information multiple times (so as to control other qubits). Present QC tools do not offer methods to explicitly indicate those behaviors, but programmers need to come up with other tools like NumPy on their own.
Therefore, a robust PL system for QC program should capture how QC programmers conceptualize such behaviors. Research in PL for QC can benefit from HCI methods, such as analyzing patterns in programming (e.g., (Denny et al., 2012; Whalley et al., 2021)), demonstrating them using diverse use cases (e.g., (Satyanarayan et al., 2016; Kim et al., 2024)), and operating user studies (e.g., (Chen et al., 2024; Myers et al., 2004, 2016b)). Such future work will make our circuit writer and result viewer interfaces operate more robustly and efficiently as well as better match to conceptual ideas.
@@ 8.2. Standardizing Interoperable Representations
Consistent data representation methods are also necessary for building interfaces, but missed currently in QC tools. For example, authentication services like ‘Sign in via Google’ can work via standardized data structures that concretely connect layers of data bases, security services, and interfaces. While implementing the prototypes, however, we frequently experienced inconsistency in API designs and their outcome formats. A major effort in our prototyping was formatting QC data in a consistent structure so as to enable on-demand user interactions on data-heavy interfaces like the circuit viewer and machine explorer. Otherwise, ad-hoc conversion had to occur every time a user clicks an element, which would add up browser-side latency.
Future research in data representation methods can work at the inter-operation of two layers: deeper, storage-efficient methods and shallower, interface-efficient methods. First, deeper representations can support storing QC data in classical databases, efficiently sharing them among users, and minimizing latency in between-application communication. Next, shallow expressions can help interface developers be less concerned with backend-side data structures, reducing the needs for ad-hoc methods. Lastly, a suite of compilers will play a key role in making those representation methods operate with each other. For instance, data visualization designs are often expressed as concise design specs (deep), which are compiled to systematically expressed visual entities (shallow) that people see on screen. Standardizing QC representation methods at varying levels will facilitate the development of useful APIs and collaborative interfaces.
@@ 8.3. Domain-specific QC Tools
QC is a specialized computing resource, intending for domain experts to solve particular types of problems. While we prototyped multiple interfaces that included problem-oriented circuit writing and result visualization, an interesting direction can be one-stop shop interfaces for specific domain areas that select and combine different techniques. For example, a QC interface for drug discovery and chemistry could help researchers to easily set superposition for molecules, encode chemical reactions as an oracle, and show outcomes in a suitable format (e.g., expressing state vectors as formulae). Similarly, simulation problems for physics can offer an editor for physical systems (e.g., a set of ions, photons) and show outcomes as animations for how those systems develop over time. The scalability of those tools will be an important concern.
Future work in domain-specific QC tools may also need to look at the interplay between graphical and character user interfaces. While graphical interfaces can reduce the gap between computational artifacts and conceptual ideas, it can be cumbersome to manipulate a lot of object by hand (e.g., dragging a few hundred gates to encode a molecule). Instead, specifying a problem as abstract formulae can make user interaction easier than graphical manipulation. Future work could also consider the interplay between QC programming and representation methods. For instance, QC programming tools can allow for directly typing in chemical formulae, and representations methods can include data about timed evolution of a physical system.
@@ 8.4. Feasible Evaluation Methods
While our goal was to propose relatively generic techniques to provide blueprints for next-generation QC tools, each interface in our work would need a designated user study for further evaluation. Prior work has done user studies on highly narrow tasks (e.g., understanding a circuit (Ruan et al., 2024a; Wen et al., 2024), interpreting quantum states (Ruan et al., 2023b), reading error rates (Ruan et al., 2023a)) with expert users who have done research in QC. Yet, future QC interfaces should be able to accommodate those who have limited backgrounds in QC but want to explore the space for potential interdisciplinary work. Evaluation studies then should be able to address the lack of sufficient QC knowledge in a feasible way; for instance, an hour-long lecture for a single-shot study is unlikely to work. Example approaches can include a single-shot user study with a well-scoped tutorial (e.g., modularization), observing how communities around tools (e.g., Qiskit user group) change in a long term, and deploying tools as course materials.